A jar contains 18 nickels, 25 dimes, and 12 quarters. What is the probability picking three dimes, replacing each after it is drawn. Show all work
Answer by richard1234(7193)

P(1 dime) = 25/(25+18+12) = 25/55 = 5/11. The probability of picking another dime is independent, and we wish to do this three times in a row, so

P(3 dimes) = (5/11)^3 = 125/1331